OpenStudy (anonymous):

80% of what number is 40?

OpenStudy (hoblos):

50

OpenStudy (anonymous):

50

OpenStudy (anonymous):

how did u get that answer? just asking because i want to be able to figure it out on my own

OpenStudy (anonymous):

40 / .8 = 50 the slash equals division. 50 * .8 = 40

OpenStudy (lalaly):

let the number be x\[\frac{80}{100} \times x=40\]solve for x

OpenStudy (anonymous):

yeahh i'm right .8 x = 40 x = 40 * (5/4) = 10 * 5 = 50 80% of 50 (answer) is 40

OpenStudy (anonymous):

how did u get .8 from 80%? moved the decimal over twice?

OpenStudy (anonymous):

Yes that is how you convert a percent to a decimal :)